# Abdoul Kader Kaboré

### v3

@@ 4 Experimental Results
Finding 20: Reasoning-based prompting significantly enhances LLM coverage across all datasets and models. However, smaller models like Mistral 7B exhibit greater limitations even with reasoning-based prompting.
4.6 [RQ6]: Execution Outcomes and Fault-Detection Capabilities
[Experiment Goal]: While coverage indicates broad execution, it doesn’t guarantee usefulness or fault detection. This experiment evaluates the execution outcomes and fault detection potential, to determine whether LLM-generated tests can effectively contribute to quality assurance
[Experiment Design]: We analyze execution behavior across three datasets (SF110, Defects4J, CMD) and multiple prompting techniques. Tests are classified as Passed (executes successfully), Failed (compilable but fails at runtime), or Timeout (exceeds 4-second limit, matching EvoSuite). We conduct failure analysis, categorizing runtime errors (ClassNotFoundException, IllegalArgumentException, NullPointerException, etc.) to distinguish flawed test logic from genuine fault detection. To assess fault-detection capabilities, we use Pitest [^13] https://github.com/hcoles/pitest to introduce controlled faults, measuring the mutation score (percentage of killed mutants) as a quantitative measure of defect identification effectiveness.
[Experiment Results]:
Test Execution Outcomes: Pass, Fail, and Timeout Rates. Our analysis reveals significant performance variations across test generation approaches. Table 17(a) shows GPT-3.5-Turbo with notable failure rates (77.11% for Defects4J, 86.75% for SF110) while Mistral 7B and Mixtral 8x7B fail 100% of CMD tests. Table 17(b) demonstrates Chain-of-Thought prompting reduces GPT-3.5-Turbo’s failures to 64.40% on Defects4J, showing clear benefits over Zero-Shot and Few-Shot approaches. Table 17(c) reveals similar CoT benefits on SF110, while Table 17(d) highlights industrial codebases’ challenges on CMD. Timeout rates reflect test complexity differences, with EvoSuite showing the highest rates (34.47% in SF110) due to its complex execution paths, while LLM-generated tests have significantly lower timeout rates (GPT-3.5-Turbo: 6.95% in SF110, 8.22% in Defects4J). Pass rates also reveal trade-offs, with GPT-3.5-Turbo achieving moderate rates (6.31% in SF110, 14.68% in Defects4J) compared to EvoSuite’s intentionally lower rates (0.27% in SF110, 2.51% in Defects4J) which align with its fault-revealing objectives. These execution metrics reveal a crucial challenge in test generation evaluation: without a reliable test oracle, distinguishing between genuine fault detection and flawed test logic remains difficult, while higher pass rates or lower timeout rates don’t necessarily indicate superior test effectiveness.
Table 17: Summary of test execution statuses (passed, failed, timed out) for LLM-generated test suites and EvoSuite.
(a) By LLM and EvoSuite across all datasets.
Models/Tool SF110 Defects4J CMD
FAILED PASSED TIMEOUT FAILED PASSED TIMEOUT FAILED PASSED TIMEOUT
GPT 3.5-turbo 86.75% 6.31% 6.95% 77.11% 14.68% 8.22% 61.66% 36.15% 2.19%
GPT 4 82.65% 5.99% 11.36%
Mistral7B 76.57% 0.85% 22.59% 86.05% 7.16% 6.80% 100% 0% 0%
Mixtral8x7B 100% 0% 0%
EvoSuite 62.26% 0.27% 34.47% 91% 2.51% 6.49%
(b) GPT 3.5-turbo and Mistral7B on Defects4J, by Prompt.
Models FAILED (%) PASSED (%) TIMEOUT (%)
ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2
GPT 3.5-turbo 75.75 75.64 64.40 84.65 85.09 14.40 13.91 29.25 5.70 10.40 10.12 10.45 6.36 9.65 4.51
Mistral7B 83.53 82.98 90.79 87.16 85.77 10.59 16.31 1.32 3.38 4.18 5.88 0.71 7.89 9.46 10.04
(c) GPT 3.5-turbo and Mistral7B on SF110, by Prompt.
Model Name FAILED (%) PASSED (%) TIMEOUT (%)
ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2
GPT 3.5-turbo 86.63 89.55 78.84 89.70 89.01 7.67 3.52 14.56 1.94 3.85 5.71 6.92 6.60 8.36 7.14
Mistral7B 72.67 97.39 41.51 83.89 87.38 0 0 0 2.78 1.46 27.33 2.61 58.49 13.33 11.17
(d) All LLMs on CMD, by Prompt.
Models FAILED (%) PASSED (%) TIMEOUT (%)
ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2 ZSL FSL COT TOT TOT2
GPT 3.5-turbo 96.49 65.05 33.60 52.21 60.96 2.63 29.13 64.40 46.90 37.67 0.88 5.83 2.00 0.88 1.37
GPT 4 68.12 94.19 80.58 86.49 83.87 0 1.16 7.77 8.11 12.90 31.88 4.65 11.65 5.41 3.23
Mistral7B 100 100 100 0 0 0 0 0 0
Mixtral8x7B 100 100 100 100 0 0 0 0 0 0 0 0
Finding 21: Test execution metrics reveal complex trade-offs in automated test generation—high failure rates (33.60-100%) reflect challenges in distinguishing genuine faults from flawed tests, lower timeout rates may indicate simpler rather than better tests, and higher pass rates don’t necessarily demonstrate superior fault detection capability.
Failure Analysis: Error Types and Root Causes. Our analysis (Table 18 and Figure LABEL:fig:failling-test-heatmap) reveals that most LLM-generated test suites fail during execution, with three predominant error types, with the y-axis of Figure LABEL:fig:failling-test-heatmap dynamically displaying error types unique to each model and test suite configuration. ClassNotFoundException (CNFE) dominates failures in larger models (GPT-3.5-Turbo: 91.76% in Defects4J, 93.89% in SF110), while IllegalArgumentException (IAE) significantly affects smaller models like Mistral 7B (22.02% in Defects4J) and both Mistral models on CMD (50%). Manual inspection confirmed these failures stem from hallucinated dependencies and invalid API usage rather than successful bug detection. Prompting strategies create notable trade-offs—ZSL and FSL generate the highest CNFE rates (GPT-4/CMD: 98.0% with ZSL vs. 75.0% with GToT), while reasoning-based approaches reduce hallucinations but increase AssertionErrors and NullPointerExceptions. Model size directly affects error patterns, with smaller models showing context-dependent failure distributions that highlight their limited repository understanding.
Table 18: Errors from failing tests by model and dataset
Model Dataset AE CNFE FNFE IOE IAE ISE IOOBE MUE NPE NFE OOME RE
GPT 3.5-Turbo Defects4J 0.88 91.76 0.02 2.91 2.92 0.47 0.56 0.02 0.38 0.03 0 0.04
SF110 0.43 93.89 0.20 0.04 4.25 0.48 0.06 0 0.48 0.07 0.03 0.04
CMD 2.60 94.80 0 0 1.20 0 0 0 1.20 0.20 0 0.00
GPT 4 CMD 5.83 83.18 0 0 2.99 0 0 0 5.70 2.16 0 0.15
Mistral 7B Defects4J 9.59 54.93 0.00 2.62 22.02 2.80 1.99 0.21 5.49 0.34 0 0.00
SF110 3.77 59.98 0.00 0.00 19.13 7.98 0.57 0 7.46 0.93 0.00 0.17
CMD 0.00 50.00 0 0 50.00 0 0 0 0.00 0.00 0 0.00
Mixtral 8x7B CMD 0.00 50.00 0 0 50.00 0 0 0 0.00 0.00 0 0.00
  • 1
    AE: AssertionError, CNFE: ClassNotFoundException, FNFE: FileNotFoundException, IOE: IOError, IAE: IllegalArgumentException, ISE: IllegalStateException, IOOBE: IndexOutOfBoundsException, MUE: MalformedURLException, NPE: NullPointerException, NFE: NumberFormatException, OOME: OutOfMemoryError, RE: RuntimeException.
Finding 22: LLM test generation failures stem primarily from model limitations rather than successful bug detection. Three key error patterns dominate: hallucinated dependencies (CNFE, 83-94% in larger models), incorrect argument validation (IAE, 19-50% in smaller models), and assertion mismatches (higher in reasoning-based models). While reasoning-based prompting reduces hallucinations, it increases assertion errors, revealing a fundamental trade-off between dependency resolution and assertion quality that varies with model architecture.
Mutation testing. Many ”Passed” test cases targeted abstract classes, interfaces, or method skeletons lacking executable logic, resulting in 0% line coverage and 0% mutation coverage in Pitest. While these tests did not produce errors, they failed to execute meaningful code, making them ineffective for fault detection. Coverage analysis in Section 4.5 further confirms this issue, with median line coverage as low as 3.57% (GPT-3.5-Turbo on SF110) and 3.03% (Mistral 7B on SF110) (Table 14). Given that most test suites were either empty (only containing test class definitions) or targeting non-executable components, computing mutation scores per class or method would not provide relevant insights.
Finding 23: Mutation testing was largely ineffective due to the non-executable nature of LLM-generated passed tests, which often targeted abstract classes or contained empty test definitions. Code coverage results confirm their failure to interact meaningfully with the code.
4.7 [RQ7]: Test Smell detection.
@@
Test Smell Distribution Across LLMs and EvoSuite. Table 19 reveals Assertion Roulette (AR), Conditional Logic Test (CLT), and Magic Number Test (MNT) as the most prevalent smells. MNT appears in nearly 100% of test suites across models, confirming Siddiq et al. (2024a)’s observation of MNT’s ubiquity. Our AR variability (73.15% with GPT-3.5/FSL on Defects4J, dropping to 19.86% with GToT on CMD) aligns with their reported range (4.6–61.3%). EH is more prevalent in EvoSuite (93.10%) than in LLMs (10–30%), partly contradicting Siddiq et al. (2024a)’s reported absence of EH in LLM outputs. This suggests reasoning-aware prompts may trigger more robust exception modeling, consistent with Finding 22. We report absence of Default Test, Verbose Test, and Dependent Test smells, supporting Finding 21 that LLM tests tend to be structurally simpler. The increase in Empty Test (EM) rates with reasoning prompts complements Finding 17, showing structured prompting improves formatting but may lead to shallow content. While Siddiq et al. (2024a) briefly report EM smells, we expand by demonstrating how prompt complexity influences EM rates across datasets.
@@
Finding 24: LLM-generated tests show simpler structures, with fewer complexity-related smells such as GF (0.0–35.96%), RA (0.0–6.17%), and CLT (0.0–14.34%). In contrast, EvoSuite consistently produces higher test smells, notably EH (93.10%), LT (76.94–82.18%), and EA (54.67–62.95%), reflecting its aggressive search-based optimization.
Dataset-Level Trends in Test Smell Prevalence. Test smell distribution reveals dataset-specific trends. Defects4J exhibits the highest Assertion Roulette rates (73.15% GPT-3.5-Turbo/FSL), indicating challenges in generating precise assertions for real-world bugs. Our findings partially align with Siddiq et al. (2024a)’s observations, but expand the analysis by comparing SF110, Defects4J, and CMD datasets. SF110 shows elevated General Fixture (93.10% in EvoSuite) and 100% Magic Number Test, while CMD demonstrates the lowest Assertion Roulette (6.4% GPT-3.5-Turbo/CoT). The analysis reveals that LLM-generated tests tend to be stateless and isolated, lacking the inter-method or state-dependent sequences typical of EvoSuite tests. This comprehensive examination illuminates how project context and software evolution influence test smell profiles across different datasets.
@@
Impact of Prompt Engineering on Test Quality. Test smell prevalence varies across LLMs, with CoT and GToT prompts improving quality. GPT-4 (CMD) reduces Assertion Roulette to 14.56%, and GPT-3.5-Turbo achieves the lowest AR at 6.4% (CoT on CMD), yet both still show nearly 100% Magic Number Test. Mistral 7B and Mixtral 8x7B yield more smells, with Mistral 7B reaching 27.97% AR on Defects4J. EvoSuite, though competitive, exhibits 100% MNT and notable AR rates (38.24–50.78%). These findings align with Siddiq et al. (2024a)’s report of universally high MNT and variable AR. Our results extend their analysis by evaluating more models and prompts, demonstrating how reasoning-based prompting reduces AR. This prompt-model-smell interplay emphasizes that improvements depend on structured prompting strategies, not just model capabilities. MNT’s persistence confirms its status as a core limitation in current test generation methods.
@@
5 Threats to Validity
@@
External Validity. The generalizability of our findings faces several limitations. Though we used three datasets (Defects4J, SF110, and CMD), these may not fully represent real-world software complexity. CMD’s small size of just two projects restricts our observations, particularly regarding the unexpected cases where GPT-3.5-Turbo outperformed GPT-4, suggesting model performance varies significantly with dataset characteristics and prompting approaches. Further experimentation on a broader dataset is necessary to confirm our observations.
@@
Internal Validity. Several internal threats could affect our results’ reliability. LLM output randomness introduces variability in generated tests, which we mitigated by executing each prompt 30 times per test suite and performing statistical analyses. Similarly, we ran 30 iterations to address EvoSuite’s variability from its evolutionary search process.
@@
Our mutation testing analysis using Pitest faced limitations as only ”Passed” tests were analyzed. Most Passed tests targeted interfaces, abstract classes, or contained empty methods, leading to 0% line coverage and mutation score. This exposes the ineffectiveness of Passed LLM-generated tests for fault detection rather than indicating a failure of mutation testing methodology. The small dataset of Passed tests further limited our findings’ generalizability.

### v4

@@ 4 Experimental Results
Finding 23: Reasoning-based prompting significantly enhances LLM coverage across all datasets and models. However, smaller models like Mistral 7B exhibit greater limitations even with reasoning-based prompting.
Relation to prior work.
Coverage is central in prior LLM-for-testing evaluations. Siddiq et al. (2024a) report high HumanEval coverage but extremely low SF110 coverage ($2\%$), with gains contingent on strong post-fix heuristics. Tang et al. (2024) find ChatGPT achieves $\approx 55\%$ statement coverage on a DynaMOSA-derived benchmark, trailing EvoSuite’s $\approx 74\%$ (30 runs). Yang et al. (2024) confirm GPT-4 lags behind EvoSuite on Defects4J, arguing hallucination-driven invalidity is the primary barrier; moreover, directly adapting In-Context Learning (ICL) techniques (CoT/RAG) from other tasks does not improve effectiveness and can reduce it for code-specialized LLMs. Yuan et al. (2024) show iterative compile-feedback (ChatTester) substantially raises compilability/executability, narrowing the gap to manually written tests.
Our results both corroborate and extend these findings. First, evaluating at the class level across Defects4J, SF110, and CMD, we show dataset characteristics strongly shape LLM coverage: relatively higher on Defects4J (potential pretraining exposure), very low on SF110 (structural complexity), and inflated on CMD (small, modular projects). Second, we systematically study prompting strategies and observe reasoning-oriented prompts (e.g., CoT/ToT/GToT) consistently improve coverage—yet still fall short of EvoSuite’s systematic exploration. This contrasts with Yang et al. (2024), where CoT and RAG—applied to code-specialized LLMs with a retrieval corpus not tailored to unit tests—are ineffective or detrimental. A plausible explanation: our prompts target developer-facing reasoning on general-purpose LLMs, whereas Yang et al. (2024) adapt In-Context Learning methods from code-generation settings to code-specialized models. Overall, our results indicate coverage improvements from ICL are model- and prompt-design dependent: effective when reasoning aligns with test-generation needs, ineffective when naively ported.
Summary of RQ5: EvoSuite consistently outperforms LLMs in structural coverage, achieving near-complete line, instruction, and method coverage. LLMs, however, display higher variability: while Defects4J shows competitive coverage (up to 82% line coverage with GPT-3.5), SF110 highlights their limitations ($5\%$) due to architectural complexity. CMD yields artificially high scores, reflecting dataset modularity rather than genuine reasoning ability. Reasoning-based prompting (CoT, ToT, GToT) narrows the gap by substantially improving LLM coverage, but systematic exploration remains a unique strength of SBST approaches like EvoSuite.
4.7 [RQ6]: Test Smell detection.
@@
Test Smell Distribution Across LLMs and EvoSuite.
Table 19 shows that Assertion Roulette (AR), Conditional Logic Test (CLT), and Magic Number Test (MNT) are the most prevalent smells in our setting. The near-ubiquity of MNT across models is consistent with Siddiq et al. (2024a), who analyze multiple smells over two benchmarks (HumanEval and SF110). Our AR variability (e.g., 73.15% with GPT-3.5/FSL on Defects4J vs. 19.86% with GToT on CMD) also falls within the ranges they report. For Exception Handling (EH), we observe higher rates in EvoSuite than in LLM outputs; this differs from patterns noted by Siddiq et al. (2024a) and suggests that reasoning prompts, as part of our prompt engineering (CoT, ToT, GToT), can elicit more explicit exception modeling. We do not observe Default Test, Verbose Test, or Dependent Test in our corpus, consistent with the broader trend that LLM-generated tests tend to be structurally simpler. Across prompt techniques, we see two systematic effects. First, under reasoning prompts (CoT/ToT/GToT), Empty Test (EM) rates increase on CMD and Defects4J, indicating that more structured prompting may improve formatting and organization yet coincide with shallower test bodies (cf. Finding 20). Second, EH tends to increase with reasoning prompts, pointing to more explicit exception scenarios. These patterns persist across model families and benchmarks; notably, MNT remains consistently dominant, with AR and CLT next in prevalence irrespective of the prompt technique.
@@
Finding 24: Across our benchmarks, LLM-generated tests tend to exhibit simpler structures, showing lower rates of complexity-related smells such as General Fixture (GF: 0.0–35.96%), Redundant Assertion (RA: 0.0–6.17%), and Conditional Logic Test (CLT: 0.0–14.34%) compared to EvoSuite. Conversely, EvoSuite consistently shows higher rates of Exception Handling (EH: 93.10%), Lazy Test (LT: 76.94–82.18%), and Eager Test (EA: 54.67–62.95%) under our configuration, which is consistent with SBST’s emphasis on structural coverage rather than stylistic simplicity.
Dataset-Level Trends in Test Smell Prevalence.
Test smell distribution reveals dataset-specific trends. Defects4J exhibits the highest Assertion Roulette rates (73.15% GPT-3.5-Turbo/FSL), indicating challenges in generating precise assertions for real-world bugs. Our findings partially align with Siddiq et al. (2024a)’s observations, but expand the analysis by comparing SF110, Defects4J, and CMD datasets. SF110 shows elevated General Fixture (93.10% in EvoSuite) and 100% Magic Number Test, while CMD demonstrates the lowest Assertion Roulette (6.4% GPT-3.5-Turbo/CoT). The analysis reveals that LLM-generated tests tend to be stateless and isolated, lacking the inter-method or state-dependent sequences typical of EvoSuite tests. This comprehensive examination illuminates how project context and software evolution influence test smell profiles across different datasets.
@@
Impact of Prompt Engineering on Test Quality.
Test smell prevalence varies across LLMs, with CoT and GToT prompts improving quality. GPT-4 (CMD) reduces Assertion Roulette to 14.56%, and GPT-3.5-Turbo achieves the lowest AR at 6.4% (CoT on CMD), yet both still show nearly 100% Magic Number Test. Mistral 7B and Mixtral 8x7B yield more smells, with Mistral 7B reaching 27.97% AR on Defects4J. EvoSuite, though competitive, exhibits 100% MNT and notable AR rates (38.24–50.78%). These findings align with Siddiq et al. (2024a)’s report of universally high MNT and variable AR. Our results extend their analysis by evaluating more models and prompts, demonstrating how reasoning-based prompting reduces AR. This prompt-model-smell interplay emphasizes that improvements depend on structured prompting strategies, not just model capabilities. MNT’s persistence confirms its status as a core limitation in current test generation methods.
@@
Relation to prior work.
Siddiq et al. (2024a) present the first systematic analysis of test smells in LLM-generated tests, applying TsDetect on HumanEval and SF110. They report near-universal Magic Number (MNT) and Lazy Test (LT) smells, with Assertion Roulette (AR) and Eager Test (EA) also prevalent, while EvoSuite exhibited higher rates of Exception Handling (EH) and LT. Their study focuses on three models (Codex, GPT-3.5, StarCoder) and two benchmarks, providing an important initial view of how smells manifest in generated tests. Our work extends this line by covering a broader spectrum of LLM families and prompting strategies across three benchmarks (Defects4J, SF110, CMD), and by contrasting them against EvoSuite. We confirm the persistence of MNT, but show that reasoning-based prompting substantially reduces AR while EvoSuite continues to exhibit higher EH, LT, and EA, highlighting complementary strengths and weaknesses between SBST and LLM-based approaches.
Summary of RQ6: LLM-generated tests show distinct smell profiles compared to EvoSuite. Magic Number Tests remain pervasive (85–100% across all settings), but reasoning-based prompts (CoT, GToT) cut Assertion Roulette rates to as low as 6.4% (vs. up to 79.3% with FSL and 38–51% with EvoSuite). EvoSuite, in turn, exhibits high Lazy Test (77–82%) and Exception Handling (93%) smells, underscoring its coverage-oriented bias. Overall, LLMs yield structurally simpler tests with fewer stateful smells (e.g., GF $\geq 36\%$ vs. 93% in EvoSuite), and prompt design emerges as a decisive factor in mitigating assertion-related issues. These results extend Siddiq et al. (2024a) by quantifying prompt-driven shifts in smell prevalence across multiple datasets and models.
5 Discussion
This section discusses the potential threats to validity, limitations of our study, and suggestions for future work.
5.1 Threats to Validity
@@
External Validity.
The generalizability of our findings faces several limitations. Though we used three datasets (Defects4J, SF110, and CMD), these may not fully represent real-world software complexity. CMD’s small size of just two projects restricts our observations, particularly regarding the unexpected cases where GPT-3.5-Turbo outperformed GPT-4, suggesting model performance varies significantly with dataset characteristics and prompting approaches. Further experimentation on a broader dataset is necessary to confirm our observations.
@@
Internal Validity.
Several internal threats could affect our results’ reliability. LLM output randomness introduces variability in generated tests, which we mitigated by executing each prompt 30 times per test suite and performing statistical analyses. Similarly, we ran 30 iterations to address EvoSuite’s variability from its evolutionary search process.
@@
5.2 Limitations and Future Work
5.2.1 Limitations.
Our study does not assess fault detection. Specifically, we did not implement (i) regression testing on paired buggy and fixed versions (i.e., generating tests on the fixed version and executing them on the buggy counterpart, as in Shamshiri et al. (2015)), (ii) the exception-oriented fault-detection setup used in recent LLM-vs-SBST comparisons (e.g., Tang et al. (2024)), which was introduced to sidestep unreliable automatically generated assertions but inherently captures only faults that manifest as runtime exceptions, or (iii) mutation-based analyses. Extending our evaluation to fault detection requires dependable oracles at scale and datasets instrumented for paired versions or mutation loops, which are outside the scope of the present study and outlined in Future Work. We also limited our comparison to EvoSuite as a single SBST baseline to maintain a clear experimental scope focused on LLM-generated test suites. Other automated test generation tools such as fuzzers, random testing, or symbolic execution frameworks were not included. We acknowledge the value of these alternatives and consider their integration an important direction for future work.
5.2.2 Future Work.
Regression testing on buggy/fixed pairs. We will re-instrument our benchmarks to enable a principled regression protocol (generate tests on the fixed version and execute them on the buggy counterpart) so that fault-revealing capability can be assessed with a controlled oracle, following established practice on Defects4J.
Principled, mutation-guided pipelines. Rather than treating mutation score as a raw end-metric, we will adopt mutation-guided loops in which surviving mutants drive subsequent test improvements (prompt augmentation/repair), with pre-filters ensuring tests are executable and assertive before mutation. This direction is supported by MuTAP(Dakhel et al. (2024)), which demonstrates that leveraging mutants as feedback can substantially raise fault-revealing effectiveness.
Hybrid SBST + LLM with full-fledged refactoring. We will combine SBST (for structural coverage and input diversity) with LLM-based refactoring beyond identifier naming (e.g., clarifying fixtures and test scaffolding, restructuring helpers, strengthening oracles, reducing brittle mocks or magic values, and improving assertion intent) so that readability and maintainability gains translate into developer effectiveness. Deljouyi et al. (2024) already show that LLM-enhanced SBST tests can improve bug-fix performance (up to +33% bugs fixed, 20-20% time) and that semantics-preserving renaming can make generated tests as readable as developer-written tests. Our next step is to generalize these benefits to structural refactoring of test code (layout, fixtures, oracles), not only surface-level names or comments.
Controlled oracles at scale. Finally, we will explore non-LLM-derived oracles (e.g., property- or metamorphic-style checks, or contract-based assertions) to reduce ambiguity in execution outcomes and make exception-only detection unnecessary, addressing a key limitation observed in recent LLM-vs-SBST comparisons.
Extended baselines. We also plan to broaden our evaluation by incorporating other automated test generation tools such as fuzzers, random test generators, and symbolic/concolic engines. This will enable a more comprehensive assessment of LLM-based test generation in relation to diverse testing paradigms.